# supplement

### v1

@@ 4.2 Quantifying User Influence
To quantify user influence, we adopted the hg-index [22], which combines the h-index [23] and g-index [29]. This metric was originally developed to evaluate scientific productivity, where citations are used as a measure of impact. In our application, we treat reposts of a user’s posts as analogous to citations, allowing us to quantify a user’s influence in social media contexts. This approach is able to balance consistent diffusion power (captured by the h-index) and the scale of reposts (captured by the g-index, similar to degree centrality).
While network centrality measures such as degree centrality and PageRank are commonly used to analyze the influence of source spreaders in social networks [1], they may disproportionately emphasize users with a single viral post. Lü et al. [30] demonstrated through a Susceptible-Infected-Recovered (SIR) model that the h-index, despite being strongly correlated with degree centrality and coreness, more appropriately evaluates node influence. However, the h-index alone does not sufficiently account for the total number of reposts, which is important in social networks. Therefore, we adopted the hg-index as a more robust measure that considers both sustained influence (through the h-index) and engagement of individual posts (through the g-index).
The hg-index was calculated using the following algorithm:
@@
    Calculate h-index: the largest h where the h-th post has at least h reposts
@@
    Calculate g-index: the largest g where the top g posts have at least $g^{2}$ total reposts
@@
    hg-index = $\sqrt{h\cdot g}$
Importantly, the hg-index mitigates the overestimation of influence based solely on total or maximum repost count, which can be skewed by a single viral post. Instead, it emphasizes users who achieve consistent engagement over time.
Based on this metric, we classified users into six influence categories: very high (top 1%), high (top 1-5%), upper-mid (top 5-10%), mid (top 10-30%), lower-mid (top 30-50%), and low (bottom 50%). Note that these categories are exclusive of their upper thresholds, meaning for example that the high category includes users above the 1% threshold but below the 5% threshold.
@@ 4.3 Construction and Analysis of Repost Cascades
For a given repost, if there are multiple potential parent posts, we generally consider the repost with the most recent timestamp as the parent. This approach is based on our inference of how Twitter’s algorithm functioned at the time of the study. Using this method, we constructed 4,882,985 repost cascades.
@@ 4.4 Virtual Timeline
We constructed a virtual timeline from sampled data and follower-followee relationships, as illustrated in Figure 2(a). This approach is inspired by the method used by Vosoughi et al. [21]. This virtual timeline simulates a single user’s timeline, providing a realistic representation of how users encounter and interact with content when scrolling through social media feeds. Notably, in our study this virtual timeline focuses exclusively on reposts.
@@
    Include only users who made at least one repost during the period, in order to remove dormant accounts
@@
    Remove subsequent reposts by a user’s followees for posts after the user reposted
This method enabled us to analyze what information users have been exposed to and how secondary spread occurs.
The virtual timeline approach allows us to simulate users’ exposure of users to content in a way that closely mimics real-world social media interactions. By focusing on reposts, we can more effectively track information diffusion through the network.
@@ 4.5 Cascading Repost Probability
To quantify the efficiency of information spread in secondary spread, we developed the concept of CRP. Using the virtual timeline approach, we consider a post as reposted if it appears in a repost cascade, allowing us to calculate the probability of a user’s secondary spread continuing, as shown in Figure 2(b).
@@
The CRP measures the likelihood of a repost being further shared, defined as:
CRP=Number of Reposed RepostsNumber of Viewed $\text{CRP}=\frac{\text{Number of Reposed Reposts}}{\text{Number of Viewed Reposts}}$ (1)
@@
To calculate the CRP for our analysis, we aggregated view counts and repost numbers across all users’ timelines. This approach allowed us to assess the extent to which users promote information spread, taking into account both the reach of their posts and the likelihood of those posts being further shared.
@@ 4.6 Ethical Considerations
5 Acknowledgement
This work was supported by JSPS KAKENHI Grant Numbers JP22K18150, JP23K28376. Initial English language improvements were assisted by AI chat services (Anthropic’s Claude 3.5 Sonnet and OpenAI’s ChatGPT). The final English language editing was provided by Editage (www.editage.jp), for which we are grateful.

### v3

@@ 4.2 Quantifying User Influence
To quantify user influence, we adopted the hg index [22], which combines the h index [23] and g index [32]. This metric was originally developed to evaluate scientific productivity, where citations are used as a measure of impact. In our application, we treat reposts of a user’s posts as analogous to citations, allowing us to quantify a user’s influence in social media contexts. This approach is able to balance consistent diffusion power (captured by the h index) and the scale of reposts (captured by the g index, similar to degree centrality).
Although network centrality measures such as degree centrality and PageRank are commonly used to analyze the influence of source spreaders in social networks [1], they may disproportionately emphasize users with a single viral post. Lü et al. [33] demonstrated through a susceptible-infected-recovered (SIR) model that the h index, despite being strongly correlated with degree centrality and coreness, more appropriately evaluates node influence. However, the h index alone does not sufficiently account for the total number of reposts, which is important in social networks. Therefore, we adopted the hg index as a more robust measure that considers both the sustained influence (through the h index) and the engagement of individual posts (through the g index).
The hg index was calculated via the following algorithm:
@@
    Calculate the h index: the largest h where the h-th post has at least h reposts
@@
    Calculate the g index: the largest g where the top g posts have at least $g^{2}$ total reposts
@@
    hg index = $\sqrt{h\cdot g}$
Importantly, the hg index mitigates the overestimation of influence on the basis solely of the total or maximum repost count, which can be skewed by a single viral post. Instead, it emphasizes users who achieve consistent engagement over time.
On the basis of this metric, we classified users into six influence categories: very high (top 1%), high (top 1–5%), upper-mid (top 5–10%), mid (top 10–30%), lower-mid (top 30–50%), and low (bottom 50%). Note that these categories are exclusive of their upper thresholds, meaning, for example, that the high category includes users above the 1% threshold but below the 5% threshold.
@@ 4.3 Construction and Analysis of Repost Cascades
For a given repost, if there are multiple potential parent posts, we generally consider the repost with the most recent timestamp as the parent. This approach is based on our inference of how Twitter’s algorithm functioned at the time of the study. However, owing to the limitations of the X (formerly Twitter) API (which does not directly provide “who reposted from whom”) and the inability to track private accounts, it is not possible to precisely reconstruct the cascade. We therefore adopt this approximation method to infer the repost chain as closely as possible. Using this method, we constructed 4,882,985 repost cascades.
@@ 4.4 Virtual Timeline
We constructed a virtual timeline from the sampled data and follower‒followee relationships, as illustrated in Figure 2(a). This approach is inspired by the method used by Vosoughi et al. [21]. This virtual timeline simulates a single user’s timeline, providing a realistic representation of how users encounter and interact with content when scrolling through social media feeds. Notably, in our study, this virtual timeline focuses exclusively on reposts.
@@
    Include only users who made at least one repost during the period to remove dormant accounts
@@
    Remove subsequent reposts by a user’s followees for posts after the user reposts
This method enables us to analyze what information users have been exposed to and how secondary spread occurs.
The virtual timeline approach allows us to simulate users’ exposure to content in a way that closely mimics real-world social media interactions. By focusing on reposts, we can more effectively track information diffusion through the network.
@@ 4.5 Cascading Repost Probability
To quantify the efficiency of information spread in secondary spread, we developed the concept of the CRP.
First, we reconstructed each user’s timeline (virtual timeline) by considering the number of times they reposted and their follower relationships to estimate which reposts were visible to each user at a given time. Next, we traced whether a viewed repost leads to a subsequent repost (further diffusion) by modeling repost-based diffusion as an information cascade [18, 21]. This approach allows us to count how many times a repost is viewed (i.e., appears on users’ timelines) and how many reposts are further shared. We then calculated, for each user, the total number of times their reposts were viewed (viewed reposts), and the total number of times those viewed reposts were further shared (reposted reposts), obtaining a per-user CRP. Finally, we grouped users by their influence categories and aggregated the counts within each group, summing the total number of viewed reposts and reposted reposts contributed by all users in that group. Hence, we obtained a single CRP value for each group that captures how likely it is that reposts from that group are to be further shared, reflecting their overall effectiveness in secondary diffusion.
Using the virtual timeline approach, we consider a post as reposted if it appears in a repost cascade, allowing us to calculate the probability of a user’s secondary spread continuing, as shown in Figure 2(b).
@@
The CRP measures the likelihood of a repost being further shared and is defined as follows:
CRP=Number of Reposted RepostsNumber of Viewed $\text{CRP}=\frac{\text{Number of Reposted Reposts}}{\text{Number of Viewed Reposts}}$ (1)
@@
This approach allowed us to assess the extent to which users promote information spread, considering both the reach of their posts and the likelihood of those posts being further shared.
4.6 Mixed Effects Logistic Regression
We employed a mixed-effects logistic regression model to determine whether user influence is a significant driver of secondary reposts after accounting for both post- and user-level heterogeneity. In this analysis, we introduced random intercepts for the source_tweet_id (i.e., the original tweet) and the topic of the user profiles (user_topic) learned by the biterm topic model[34]. This approach helps control for unobserved differences in content-specific appeal and individual interests while allowing us to assess the influence of specific predictors on the likelihood of a repost.
We used the lme4 package [35] in R to fit the model as follows:
glmer(
  is_retweeted ~ followee_influence + repost_hour + sender_followers_count +
  user_followers_count + (1 | user_topic) + (1 | source_tweet_id),
  data = data,
  family = binomial(link = "logit"),
  nAGQ = 0,
  control = glmerControl(optimizer = "bobyqa", optCtrl = list(maxfun = 1e+05))
Ψ
The binary outcome variable, is_retweeted, indicates whether the repost, which is shown to a user, triggered an additional repost by that user. As fixed effects, the model incorporates (1) the user’s influence category (sender_influence), coded via intercept (reference) coding with low as the baseline; (2) the time of day the repost was shown (repost_hour), coded with sum contrasts to capture diurnal patterns, which were referenced against Japan’s National Statistics on media use [36]; (3) the sender’s follower count (sender_followers_count); and (4) the viewer’s follower count (user_followers_count). For both follower-count variables, we add 1 before applying the log transformation and then standardize the resulting values to facilitate model convergence and interpretation.
To address the computational challenges arising from a vast dataset, we reduced the sample by collecting approximately 0.1% of all positive (further reposted) cases and randomly drawing negative cases (displayed but not reposted) at twice the frequency of the positives. For comparability, negative instances were drawn only from tweets with a source_tweet_id present in the positive set and from timelines of users who appeared in the positive set. Because this sampling procedure artificially alters the original ratio of positive to negative outcomes, the model’s intercept and coefficients cannot be interpreted as absolute probabilities of reposting. Nonetheless, they effectively capture relative differences, such as contrasts among influence categories.
Although a user-level random intercept such as user_id would ordinarily capture individual differences, the data were too sparse at the single-user level—most users only reposted the same tweet once. Instead, we applied a random intercept for the topic of the user profiles to stabilize the estimation and account for higher-level differences reflecting user interests. The Supplementary Material S2 details this topic of user profiles. We also added a random intercept for source_tweet_id to control for postspecific characteristics influencing repost likelihood, such as content, emotional tone, or structural features. We used a binomial error distribution with a logit link function, specified nAGQ = 0, and employed the bobyqa optimizer to manage the computational load.
This model clarified the role of user influence in secondary spreading by integrating post and user random effects and controlling for key predictors such as the hour of the day, follower count, and influence category.
4.7 Max Depth and Structural Virality of First-Reposted Users
This analysis selected posts with more than 5,000 reposts to capture instances of substantial online diffusion. For each post, the first user who reshared it (the first-reposted user) was identified, and the corresponding repost cascade was isolated by tracing all subsequent shares. Users were categorized by influence level—ranging from low to very high—on the basis of criteria such as follower counts, previous engagement metrics, and other indicators of potential reach.
To quantify the complexity and depth of the repost cascades, the first was structural virality, which, following the approach of Goel et al. [28] and Vosoughi et al. [21], was computed as the average shortest path length among all pairs of users in a given cascade. Higher values of structural virality indicate multistep or chain-like diffusion, whereas lower values indicate broadcast-like patterns with fewer resharing steps. The second measure was the maximum depth, which was calculated by determining the longest path from the first reposted user to the final user in the chain, thus indicating how many layers of sharing emerged from the initial repost.
These metrics were then compared across influence categories to assess whether highly influential users consistently spurred more extensive diffusion. By evaluating structural virality and maximum depth in the context of the initial user influence level, this study elucidated how early engagement by users with high influence can catalyze further resharing and produce cascading viral chains throughout the platform.
@@ 4.8 Ethical Considerations
4.9 Limitations
  • Language Constraints
    The primary dataset used in this study consists of Japanese-language content. Although the English data we used cover a single day’s worth of posts, this is insufficient for thorough experimentation. Furthermore, English data were not originally collected for this experiment. limiting its applicability to the current research goals. Consequently, comprehensive experiments and analyses involving English and other languages remain as future work.
  • Sampling Bias
    The data analyzed in this study were collected via the X (formerly Twitter) API, meaning it represents a sampled subset rather than the platform’s entire dataset. Moreover, private accounts are excluded from data collection, which may introduce a certain degree of sampling bias.
  • Timeline Assumption
    This study assumes that the timeline on X is chronological. Although we briefly discuss advertisements and algorithm-driven content reordering in the Supplementary Material S5, This consideration is not comprehensive.
These limitations should be considered when interpreting this study’s findings and conclusions. Future work will aim to address these issues by expanding linguistic coverage, employ broader data collection methods, and more thoroughly examining the impact of algorithmic content ordering.
5 Acknowledgments
This work was supported by JSPS KAKENHI Grant Numbers JP22K18150 and JP23K28376. The English editing was provided by Editage (www.editage.jp), for which we are grateful.
6 Additional Information
6.1 Competing interests
The authors declare that they have no competing interests.
7 Data and Code availability
All datasets and codes used in this work are available at https://github.com/asahi-research/secondary_spreads_of_influencers.